Electrons are accelerated by 2150V in an electron microscope.

What is the maximum possible resolution?

(Answer in meters)

Please show all work.
Answer #1

We know

\lambda =\frac{h}{p}      (De-Brogle equation) p is momentum and h is plank constant.

we also have P=\sqrt{2mE}    ( here m is mass and E is energy )

from above both eqution      

\lambda =\frac{h}{\sqrt{2mE}}     ---------------------------------------------------(1)

also E=qV ( q charge and V is voltage)

putting value of E in eq 1

\lambda =\frac{h}{\sqrt{2mqV}}     (This is the equation for maximum resolution)

putting values of h=6.62\times 10^{^{-34}}, m=9.1\times10-31kg, q=1.6\times10-19 c, V=2150v

\lambda =\frac{6.62\times 10^{-34}}{\sqrt{2\times 9.1\times 10^{-31}\times 1.6\times 10^{-19}\times 2150}}      =2.64\times1014 meter (ans)